Wylie’s Watering Hole sells fruit smoothies. Last year, Wylie sold a total of 24,000 fruit smoothies. Pineapple smoothies outsold Coconut smoothies 2 to 1. Sales of Mango smoothies were the same as sales of pineapple. Fixed costs for the company are $16,168 and additional information follows:

Product Unit Sales
Prices
Unit Variable
Cost
Pineapple $2.60 $0.90
Coconut $3.85 $1.45
Mango $3.30 $1.50

a. Calculate the sales mix percentage of all three products (based upon the number of smoothies).

Pineapple Coconut Mango
Sales mix (%) Answer Answer Answer

b. Calculate the weighted average contribution margin for all three types of drinks.

Pineapple Coconut Mango Total
Wtd. avg. CM Answer Answer Answer Answer

c. Calculate the break-even volume in number of total smoothies.

Answer

d. Calculate the break-even point in revenue.

Round answer to the nearest dollar.

$Answer
Answer #1

a. Calculate the sales mix percentage of all three products (based upon the number of smoothies).

Pineapple Coconut Mango
Sales mix (%) 2/5 = 40% 20% 40%

b. Calculate the weighted average contribution margin for all three types of drinks.

Pineapple Coconut Mango Total
Wtd. avg. CM 1.7*40% = 0.68 2.4*20% = 0.48 1.8*40% = 0.72 1.88

c. Calculate the break-even volume in number of total smoothies.

16168/1.88 = 8600 Units

d. Calculate the break-even point in revenue.

Round answer to the nearest dollar.

8600*3.13 = 26918